# Emergent vorticity asymmetry of one and two-layer shallow water system captured by a next-order balanced model

### v1

@@ 2.2.1 Set-ups of the simulations
The simulations are set in an idealized mid-latitude ocean mixed-layer modeled as a one-and-a-half-layer shallow water system. It is translated to our dimensional parameters of height $H=100\text{ m}$, Coriolis parameter $f=10^{-4}\text{ s}^{-1}$, and a reduced gravity constant $g=1\text{ m/s}^{2}$ (with the usual prime omitted to match the notation of the rest of the section). These parameters imply a deformation radius $\sqrt{gH}/f$ of 100 km. The domain is doubly-periodic with $Lx=Ly=12\pi\times 100$ km. All the freely decaying simulations have the Burger number (3) equal to one, that is we nondimensionalize with $L=100$ km and the nondimensional domain is $12\pi\times 12\pi$. A typical flow speed at energetic regions of the upper ocean (e.g.: the Gulf Stream) is around $U=1.2$ m/s. This gives a Rossby number (2) of $\varepsilon=0.12$ while a more quiescent region has $U=0.1$ m/s and $\varepsilon=0.01$. Our simulation will explore this range of realistic Rossby numbers. Of course, the power of nondimensionalization allows our simulation to apply to the atmosphere as well. We do not explore higher Rossby numbers since too high a Rossby number while keeping the Burger number fixed will result in $\{\varepsilon/Bu\}h/H>1$, a drying of part of the domain. This is unrealistic for large-scale ocean and atmosphere applications and cannot be simulated using our pseudospectral code.
@@
Vorticity asymmetry emerges from the evolution of the shallow water model. We evolve the simulation pseudospectrally using Dedalus (Burns et al., 2020) for 500 nondimensional time. The horizontal resolution is $1024\times 1024$ Fourier modes. We dissipate the small-scale for numerical stability using a fourth-order hyper-dissipation
$\nu\nabla^{4}(\cdot)\quad\mbox{with}\quad\nu=5.5\times 10^{-6}.$ (30)
@@ 2.2.2 Simulation results
Figure 1 shows the comparison of snapshots of vorticity and height for the two models at $t/T=200$ with $\varepsilon=0.1$. The domains are filled with roaming vortices with a bias towards anti-cyclonic ones. The dominant length scale becomes large compared to the initial condition, which is associated with EKE conversion to APE (cf. Figure 4) and the inverse cascade.
@@
Ensemble statistics is appropriate for studying the emergent vorticity asymmetry of the shallow water model and whether the SWQG+1 model can capture it. The left of Figure 2 shows the time evolution of vorticity skewness of the size ten ensemble of simulations of the shallow water model and SWQG+1, for $\varepsilon=0.1$. The initial conditions are constructed to have symmetric vorticity and indeed the vorticity skewness starts at zero for all simulations. All simulations’ vorticity skewness trends to negative until they reach statistical equilibrium. The spread between members of the ensembles becomes larger over time, expected for the chaotic dynamics of shallow water turbulence. However, the ensemble means of vorticity skewness of the two models agree remarkably well during the entire time series. They lie within $1/\sqrt{10}$ times the ensemble standard deviation from each other for most of the time series. We use this range since it is the standard scaling of error of a Monte-Carlo estimate of the ensemble mean. The close agreement indicates strongly that SWQG+1 model can capture the mechanism that leads to vorticity asymmetry in the shallow water model (Arai & Yamagata, 1994; Graves et al., 2006).
Figure 2: Left: time series of vorticity ($\zeta$) skewness for the $\varepsilon=0.1$ simulations from the shallow water model as well as SWQG+1. The lighter lines are the individual ensemble members. The darker lines are the ensemble mean and the $1/\sqrt{10}$ of the ensemble standard deviation is the filled color around the mean. Right: the vorticity ($\zeta$) skewness at $t/T=200$ for $\varepsilon=0.01,0.03,0.05,0.07,0.1,0.12$. The error bar is $1/\sqrt{10}$ of the ensemble standard deviation.
The agreement extends to all Rossby numbers explored. The right of Figure 2 shows the ensemble mean and $1/\sqrt{10}$ times the ensemble standard deviation of vorticity skewness for the shallow water model and SWQG+1. We show the data at nondimensional time $t/T=200$. We see that vorticity skewness becomes more negative as the Rossby number gets larger. Polvani et al. (1994, FIG. 10) shows a similar trend except with the Froude number, which is equal to the Rossby number for all our simulations where the Burger number is one.
@@
which is equal to Froude number squared. This is consistent with the fact that vorticity skewness is proportional to the Froude number (Polvani et al., 1994).
@@
The freely evolving simulation in the shallow water model is a simple setting where we can investigate the properties of the SWQG+1 model. We have not been able to show that the SWQG+1 model conserves energy. Here instead we diagnose the energy and show that it is well-behaved. The left of Figure 4 shows the time series of the total energy as well as the EKE and APE components (9) for the evolution of the $\varepsilon=0.1$ ensemble. The EKE decreases as it is converted to the APE, a signature of the inverse cascade. The total energy is well-conserved such that more than 90% of it is in the final state. However, it is noticeable that the SWQG+1 loses more energy, compared to the shallow water model, even though we have used the same small-scale dissipation operator. The shallow water model dissipates the physical variables separately, while the SWQG+1 model dissipates the PV. Dissipation in the shallow water model does not translate to dissipation of the PV, because of the nonlinearity in the shallow water PV (6). This is well-known for the Ertel PV of the Boussinesq system, where the dissipation of the physical variables translates to a source term in the PV equations (Haynes & McIntyre, 1987). We will not explore the possibility of including source terms in the PV equation in this work any further. The total energy is minimally dissipated and the dissipation effects are not significant.
It is important to remark that the total energy of the SWQG+1 model is not monotonic with time. It does not conserve this definition of energy, though overall shallow water total energy characterizes the evolution of the model well. In contrast, the total energy at the QG-level (22) does not behave well. Figure 5 shows the time series of the QG-level total energy. They do not even decrease overall and have a much larger ensemble spread.
@@
We have commented in Section 2.1.2 that total energy at the QG level is proportional to $C_{q}$ (21). It is the SWQG+1 representation of the mean PV ($\left\langle{q}\right\rangle$) of the shallow water model. Figure 5 compares the two. Overall, the two values are similar. This confirms the assumption made in the derivation that the difference will be of $O(\varepsilon^{2})$. Mean PV evolves in shallow water due to the correlation between PV and divergence:
@@
Therefore $C_{q}$ is the only representation of the evolution of the mean of PV ($\left\langle{q}\right\rangle$) in the SWQG+1 model.
@@ 3 Multi-layer shallow water
We extend the QG+1 model to a multi-layer configuration. Here we only present the two-layer version, but the extension to more layers is straightforward following the steps laid out here. A simulation of the two-layer version of SWQG+1 under baroclinic instability shows that it can capture the vortex asymmetry in nonlinear baroclinic waves.
@@ 3.1 The two-layer shallow water system and the derivation of the two-layer SWQG+1 model
We start with the two-layer shallow water system on the $f$-plane with two equal layers of height $H$ and inter-layer reduced gravity $g^{\prime}$
@@
All nondimensional numbers are the same as the one for the one-layer shallow water, except we use the baroclinic Burger number
@@
The operator on the left-hand side will appear in all higher-order inversions and we name it $\mathcal{L}$ for convenience.
@@ 3.2.1 Set-ups of the simulations
The instability is triggered by the seeding of small magnitude noise, to height in the shallow water simulation, and to PV in the SWQG+1 simulation. We again use Dedalus to simulate the nonlinear evolution. The resolution is $Nx\times Ny=512\times 256$ modes. Everything else about the numerics are the same as the simulation in Section 2.2 except the fourth-order hyper-dissipation constant is larger
@@ 3.2.2 Simulation results
The SWQG+1 captures the vorticity evolution of the baroclinic life cycle to remarkable accuracy. Figure 7 displays snapshots of the barotropic ($(\zeta_{1}+\zeta_{2})/2$) and baroclinic ($(\zeta_{1}-\zeta_{2})/2$) vorticity (in color) and divergence (in contour) during the growth of the baroclinic instability. Note that the time of the snapshot is offset. This is because the growth of the instability from small-scale noise happens at different times in the model, due to the chaotic nature of the instability. The match is indeed impressive. Not only does the SWQG+1 model capture the asymmetric evolution of the cyclonic and anti-cyclonic part of the baroclinic wave, it is also able to recover the associated divergence fields. The divergence is of crucial atmospheric interest as a convergent region in the lower layer corresponds to precipitation. For our example, baroclinic convergence correlates with baroclinic cyclonic vorticity, and baroclinic divergence is large near regions of barotropic strain. On the other hand, barotropic divergence is uniformly small. This is reminiscent of strain-induced fronts configuration studied in Hoskins & Bretherton (1972). We will explore further strain-induced fronts in Section 3.3.
@@
Figure 7: The vorticity field of the evolution of the jets for the shallow water model (left) and SWQG+1 model (right). The contour is divergence $\{\varepsilon\}\delta/f$ at $[-0.01,-0.005,0.005,0.1]$, with the solid being positive. The shallow water snapshot is taken at $t/T=30$ while the SWQG+1 snapshot is taken at $t/T=24$.
@@
The match of the vorticity snapshots translates to the ability of the SWQG+1 to capture the vorticity skewness of the simulation. Figure 8 shows the agreement of the overall trend extending to a large time ($t/T=100$), for both layers. The vorticity is now skewed towards positive, in comparison to the freely decaying result. This change in vorticity skewness is due to the new effects of the baroclinic mode now included in the two-layer model. The vortex stretching due to the correlation of baroclinic vorticity and divergence drives the initial growth of positive vorticity skewness shown in Figure 8. This effect is entirely missing in the one-layer simulation.
@@
Our results agree with the ones of Lambaerts et al. (2012) in the upper layer for the onset of instability. However, their lower layer does not have a mean jet while ours does thus the results differ.
@@
Figure 8: Vorticity skewness of two-layer shallow water (left) and SWQG+1 (right) during the evolution of unstable jets simulations.
@@ 3.3 Baroclinic strain-driven fronts modeled by SWQG+1
The two-layer simulation reveals strain-driven fronts are important in the dynamics and contribute to the positive vorticity skewness. Here we study a minimal model of strain-driven fronts by adapting the set-up of Hoskins & Bretherton (1972) to the two-layer.
We study a strain cold filament across the $y$-direction. That is, we ignore the $x$-variation for all perturbation fields. The filament has a center with a large $h_{2}$ perturbation. We prescribe the perturbation PV field to be Gaussian
@@
Note that the operator $\mathcal{L}$ now only has $y$-derivatives. Without the strain field ($\alpha=0$), there would be no response in $v$ since $G=0$. This is expected since the filament is a geostrophically balanced solution to the shallow water equations.
@@
We solve the system with $Bu=1$ and $Ro=0.3$. Figure 9 shows the divergence
@@
overlaid on the height perturbations. In the top layer, a cold filament has cyclonic vorticity, which is correlated with convergence. This correlation is a key feature of fronts and explains the cyclonic bias of vorticity in the two-layer simulations by appealing to vortex stretching. However, layered shallow water is different from the surface buoyancy gradient driven Boussinesq frontogenesis of Hoskins & Bretherton (1972) in that this convergence will smooth away the height perturbation and release the stored APE. Therefore shallow water fronts would be less strong with no finite time singularity, compared to the Boussinesq case.

### v3

@@ 2.2.1 Set-ups of the simulations
The simulations are set in an idealized mid-latitude ocean mixed-layer modeled as a one-and-a-half-layer shallow water system. It is translated to our dimensional parameters of height $H=100\text{ m}$, Coriolis parameter $f=10^{-4}\text{ s}^{-1}$, and a reduced gravity constant $g=1\text{ m/s}^{2}$ (with the usual prime omitted to match the notation of the rest of the section). These parameters imply a deformation radius $\sqrt{gH}/f$ of 100 km. The domain is doubly-periodic with $Lx=Ly=12\pi\times 100$ km. All the freely decaying simulations have the Burger number (3) equal to one, that is, we nondimensionalize with $L=100$ km and the nondimensional domain is $12\pi\times 12\pi$. A typical flow speed at energetic regions of the upper ocean (e.g., the Gulf Stream) is around $U=1.2$ m/s. This gives a Rossby number (2) of $\varepsilon=0.12$ while a more quiescent region has $U=0.1$ m/s and $\varepsilon=0.01$. Our simulation will explore this range of realistic Rossby numbers. Of course, the power of nondimensionalization allows our simulation to apply to the atmosphere as well. We do not explore higher Rossby numbers since too high a Rossby number while keeping the Burger number fixed will result in $\{\varepsilon/Bu\}h/H>1$, a drying of part of the domain. This is unrealistic for large-scale ocean and atmosphere applications and cannot be simulated using our pseudospectral code.
@@
Vorticity asymmetry emerges from the evolution of the shallow water model. We evolve the simulation pseudospectrally using Dedalus (Burns et al., 2020) for 500 nondimensional time. The horizontal resolution is $1024\times 1024$ Fourier modes. We dissipate the small-scale for numerical stability using a fourth-order hyper-dissipation just large enough to absorb small-scale noise, as diagnosed by the vorticity spectrum
$\nu\nabla^{4}(\cdot)\quad\mbox{with {\color[rgb]{0,0,0}non-dimensional value}}\quad\nu=5.5\times 10^{-6}.$ (32)
@@ 2.2.2 Simulation results
Figure 1 shows the comparison of snapshots of vorticity and height for the two models at $t/T=200$ with $\varepsilon=0.1$. The domains are filled with roaming vortices with a bias towards anticyclonic ones. The dominant length scale becomes large compared to the initial condition, which is associated with EKE conversion to APE (cf. Figure 4) and the inverse cascade. For the one-layer simulations, balanced evolution of vortices does not generate motions with strong divergence, and its effect on the dynamics is minimal. Therefore, we do not further investigate divergence for the one-layer simulations.
@@
Ensemble statistics are appropriate for studying the emergent vorticity asymmetry of the shallow water model and whether the SWQG+1 model can capture it. The left of Figure 2 shows the time evolution of vorticity skewness of the size ten ensemble of simulations of the shallow water model and SWQG+1, for $\varepsilon=0.1$. The initial conditions are constructed to have symmetric vorticity, and indeed, the vorticity skewness starts at zero for all simulations. All simulations’ vorticity skewness trends towards negative. The spread between members of the ensembles becomes larger over time, as expected for the chaotic dynamics of shallow water turbulence. However, the ensemble means of vorticity skewness of the two models agree remarkably well during the entire time series. They lie within $1/\sqrt{10}$ times the ensemble standard deviation from each other for most of the time series. We use this range since it is the standard scaling of error of a Monte-Carlo estimate of the ensemble mean. The close agreement indicates strongly that SWQG+1 model can capture the mechanism that leads to vorticity asymmetry in the shallow water model (Arai & Yamagata, 1994; Graves et al., 2006).
Figure 2: Left: time series of vorticity ($\zeta$) skewness for the $\varepsilon=0.1$ simulations from the shallow water model as well as SWQG+1. The lighter lines are the individual ensemble members. The darker lines are the ensemble mean, and the $1/\sqrt{10}$ of the ensemble standard deviation is the filled color around the mean. Right: the vorticity ($\zeta$) skewness at $t/T=200$ for $\varepsilon=0.01,0.03,0.05,0.07,0.1,0.12$. The error bar is $1/\sqrt{10}$ of the ensemble standard deviation.
The agreement extends to all Rossby numbers explored. The right of Figure 2 shows the ensemble mean and $1/\sqrt{10}$ times the ensemble standard deviation of vorticity skewness for the shallow water model and SWQG+1 at nondimensional time $t/T=200$. We see that vorticity skewness becomes more negative as the Rossby number gets larger. Polvani et al. (1994, FIG. 10) shows a similar trend except with the Froude number, which is equal to the Rossby number for all our simulations where the Burger number is one.
@@
which is equal to the Froude number squared. This is consistent with the fact that vorticity skewness is proportional to the Froude number (Polvani et al., 1994).
@@
The freely evolving simulation in the shallow water model is a simple setting where we can investigate the properties of the SWQG+1 model. We have not been able to show that the SWQG+1 model conserves energy. Instead, we diagnose the energy and show that it mimics the behavior of the energy of the shallow water model. The left of Figure 4 shows the time series of the total energy as well as the EKE and APE components (9) for the evolution of the $\varepsilon=0.1$ ensemble. The EKE decreases as it is converted to the APE, a signature of the inverse cascade. The total energy is well-conserved such that more than 90% of it is in the final state. The SWQG+1 total energy evolution closely follows the one for the full shallow water model, and this is true for all Rossby numbers explored (not shown). However, it is noticeable that the SWQG+1 loses more energy, compared to the shallow water model. The shallow water model dissipates the physical variables separately, while the SWQG+1 model dissipates the PV. Dissipation in the shallow water model does not translate to dissipation of the PV, because of the nonlinearity in the shallow water PV (6). This is well-known for the Ertel PV of the Boussinesq system, where the dissipation of the physical variables translates to a source term in the PV equations (Haynes & McIntyre, 1987). We will not explore the possibility of including source terms in the PV equation in this work any further. The total energy is minimally dissipated, and the dissipation effects are not significant.
It is important to remark that the total energy of the SWQG+1 model is not monotonic with time, though this is hard to notice in the timeseries in Figure 4. This implies that the SWQG+1 model does not conserve the shallow water energy, at least with this choice of dissipation of PV. An alternative reasonable measurement of energy is the total energy at the QG-level (22). But it does not even decrease overall and has a much larger ensemble spread, as shown by the time series in Figure 5. This further shows our simulated regime is beyond simple QG dynamics.
@@
We have commented in Section 2.1.2 that for SWQG+1 total energy at the QG level (22) is proportional to $C_{q}$ (21). It is the SWQG+1 representation of the mean PV ($\left\langle{q}\right\rangle$) of the shallow water model. Figure 5 compares the two. Overall, the two values are similar. This confirms the assumption made in the derivation that the difference will be of $O(\varepsilon^{2})$. Mean PV evolves in shallow water due to the correlation between PV and divergence:
@@
Therefore, $C_{q}$ is the only representation of the evolution of the mean of PV ($\left\langle{q}\right\rangle$) in the SWQG+1 model.
@@ 3 Multi-layer shallow water
We extend the QG+1 model to a multi-layer configuration. Here we only present the two-layer version, but the extension to more layers is straightforward, following the steps laid out here. A simulation of the two-layer version of SWQG+1 under baroclinic instability shows that it can capture the vortex asymmetry in nonlinear baroclinic waves.
@@ 3.1 The two-layer shallow water system and the derivation of the two-layer SWQG+1 model
We start with the two-layer shallow water system on the $\beta$-plane with two equal layers of height $H$ and inter-layer reduced gravity $g^{\prime}$
@@
All nondimensional numbers are the same as those for the one-layer shallow water, except we use the baroclinic Burger number
@@
The operator on the left-hand side will appear in all higher-order inversions, and we name it $\mathcal{L}$ for convenience.
@@ 3.2.1 Set-ups of the simulations
The instability is triggered by the seeding of small magnitude noise, to height in the shallow water simulation, and to PV in the SWQG+1 simulation. We again use Dedalus to simulate the nonlinear evolution. The resolution is $Nx\times Ny=512\times 256$ modes. Everything else about the numerics are the same as the simulation in Section 2.2 except the fourth-order hyper-dissipation constant is larger with non-dimensional value
@@ 3.2.2 Simulation results
The SWQG+1 captures the vorticity evolution of the baroclinic life cycle to remarkable accuracy. Figure 7 displays snapshots of the barotropic ($(\zeta_{1}+\zeta_{2})/2$) and baroclinic ($(\zeta_{1}-\zeta_{2})/2$) vorticity (in color) and divergence (in contour) during the growth of the baroclinic instability. Note that the time of the snapshot is offset. This is because the growth of the instability from small-scale noise happens at different times in the model, due to the chaotic nature of the instability. However, the growth rate is remarkably similar (see Figure 8). The match is indeed impressive. Not only does the SWQG+1 model capture the asymmetric evolution of the cyclonic and anticyclonic part of the baroclinic wave, it is also able to recover the associated divergence fields. Note that one could infer the same divergence, using the $\omega$-equation. The difference of SWQG+1 is that the full velocity participates in the advection, while in SWQG, the higher-order divergence is implicit and is not used in the advection of PV. The divergence is of crucial atmospheric interest as a convergent region in the lower layer corresponds to precipitation. For our example, baroclinic convergence correlates with baroclinic cyclonic vorticity, and baroclinic divergence is large near regions of barotropic strain. On the other hand, barotropic divergence is uniformly small. This is reminiscent of the strain-induced fronts configuration studied in Hoskins & Bretherton (1972). We will explore further strain-induced fronts in Section 3.3. The similarity of the evolution for the two models go beyond this early snapshot. In the supplementary materials, we provide videos of the $\{\varepsilon\}=0.1$ simulations.
@@
Figure 7: The vorticity field of the evolution of the jets for the shallow water model (left) and SWQG+1 model (right). The contour is divergence $\{\varepsilon\}\delta/f$ at $[-0.1,-0.005,0.005,0.1]$, with the solid being positive. The shallow water snapshot is taken at $t/T=30$ while the SWQG+1 snapshot is taken at $t/T=24$.
@@
The match of the vorticity snapshots translates to the ability of the SWQG+1 to capture the vorticity skewness of the simulation. Figure 8 shows the agreement of the overall trend over a large time ($t/T=100$), for both layers. The vorticity is now skewed towards positive, in comparison to the freely decaying result. This change in vorticity skewness is due to the new effects of the baroclinic mode now included in the two-layer model. The vortex stretching due to the correlation of baroclinic vorticity and divergence drives the initial growth of positive vorticity skewness shown in Figure 8. This effect is entirely missing in the one-layer simulation. The vorticity skewness is an ageostrophic phenomenon. We diagnose the first local maxima of the vorticity skewness, as it is an indicator of the strength of the ageostrophic frontogenesis. Figure 9 shows clearly that this measurement of vorticity skewness is linearly proportional to the Rossby number, and SWQG+1 is effective over a large parameter space. However, it is worth noting that this close match between shallow water and SWQG+1 might not hold across all Burger numbers. Our chosen regime has minimal spontaneous gravity wave emission, even for high Rossby number. This might not be the case for another choice of Burger number. We leave further exploration of other parameters to further study.
@@
Our results agree with those of Lambaerts et al. (2012) in the upper layer for the onset of instability. However, their lower layer does not have a mean jet while ours does, thus the results differ.
@@
Figure 8: Vorticity skewness of two-layer shallow water (left) and SWQG+1 (right) during the evolution of unstable jets simulations. The circles and crosses mark the first local maxima of vorticity skewness, which is further explored in Figure 9.
Figure 9: The first local maxima of vorticity skewness of a set of simulations with varying Rossby numbers.
@@ 3.3 Baroclinic strain-driven fronts modeled by SWQG+1
The two-layer simulation reveals that strain-driven fronts are important in the dynamics and contribute to the positive vorticity skewness, which is opposite to the results of our one-layer simulations. This is a new feature of the two-layer shallow water simulations and worth further investigation. Here we study a minimal model of strain-driven fronts by adapting the set-up of Hoskins & Bretherton (1972) to the two-layer.
We study a strain-driven cold filament across the $y$-direction. That is, we ignore the $x$-variation for all perturbation fields. The filament has a center with a large $h_{2}$ perturbation. We prescribe the perturbation PV field to be Gaussian
@@
Note that the operator $\mathcal{L}$ now only has $y$-derivatives. Without the strain field ($\alpha=0$), there would be no response since $G=0$. This is expected since the filament is a geostrophically balanced solution to the shallow water equations.
@@
We solve the system with $Bu=1$ and $\varepsilon=0.3$. Figure 10 shows the divergence
@@
overlaid on the height perturbations. In the top layer, a cold filament has cyclonic vorticity, which is correlated with convergence. This correlation is a key feature of fronts and explains the cyclonic bias of vorticity in the two-layer simulations by appealing to vortex stretching: baroclinic cyclonic vorticity is correlated with baroclinic negative divergence. Note that the above equation is equivalent to the “$\omega$-equation”. Therefore, it is implicit in the SWQG model. However, SWQG does not use this ageostrophic velocity in its advective evolution. It does not capture the vorticity skewness evolution.
Finally, a comparison with the surface buoyancy gradient driven Boussinesq frontogenesis of Hoskins & Bretherton (1972) is warranted. In the layered shallow water set-up, the convergence will smooth away the height perturbation and release the stored APE. Shallow water fronts would be less strong with no finite time singularity, compared to the Boussinesq case.